# Efficient Batch and Recursive Least Squares for Matrix Parameter Estimation with Application to Adaptive MPC

### v1

@@ IV Independent Identical Column Weighting
Propositions 5 and 6 respectively give the independent identical column weighting batch and recursive least squares methods to minimize cost function $J_{k}$ given by (IV).
@@
Let $N\geq 0$. For all $0\leq k\leq N$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let ${\Gamma}_{k}\in{\mathbb{R}}^{p\times p}$ be positive definite. Furthermore, let ${\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let ${R}\in{\mathbb{R}}^{n\times n}$ be positive definite. Then, $J_{N}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (IV), has a unique minimizer given by
$\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{N}(\hat{\theta})=A_{N}^{-1}b_{N},$ (33)
@@
$A_{N}$ $\triangleq R+\sum_{i=0}^{N}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i},$ (34)
$b_{N}$ $\triangleq R\theta_{0}+\sum_{i=0}^{N}\phi_{i}^{\rm T}\Gamma_{i}y_{i}.$ (35)
Proof.
Note that (IV) can be rewritten as (18) where, for all $k\geq 0$ and $j=1,\ldots,m$, $\Gamma_{j,k}=\Gamma_{k}$ and $R_{j}=R$. It then follows from Proposition 3 that $\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{N}(\hat{\theta})=\begin{bmatrix}A_{N}^{-1}b_{1,N}&\cdots&A_{N}^{-1}b_{m,N}\end{bmatrix}=A_{N}^{-1}\begin{bmatrix}b_{1,N}&\cdots&b_{m,N}\end{bmatrix}$, where $A_{N}\in{\mathbb{R}}^{n\times n}$ is defined in (34) and $b_{j,N}$ is defined in (25). Finally, note that $b_{N}=[b_{1,N}\ \cdots\ b_{m,N}]$, thus yielding (33). ∎
Proposition 6.
For all $k\geq 0$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let ${\Gamma}_{k}\in{\mathbb{R}}^{p\times p}$ be positive definite. Furthermore, let ${\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let ${R}\in{\mathbb{R}}^{n\times n}$ be positive definite. Then, for all $k\geq 0$, $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (IV), has a unique minimizer, denoted as
$\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta})=\theta_{k+1}.$ (36)
Moreover, for all $k\geq 0$, $\theta_{k+1}\in{\mathbb{R}}^{n\times m}$ is given by
@@
where $P_{0}\triangleq R^{-1}$ and, for all $k\geq 0$, $P_{k}\in{\mathbb{R}}^{n\times n}$ is positive definite, hence nonsingular. Finally, for all $k\geq 0$, $P_{k+1}$ can be expressed recursively as
@@
Note that (IV) can be rewritten as (18) where, for all $k\geq 0$ and $j=1,\ldots,m$, $\Gamma_{j,k}=\Gamma_{k}$ and $R_{j}=R$. It then follows from Proposition 4 that, for all $k\geq 0$,
$\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta})=\begin{bmatrix}\theta_{1,k+1}&\cdots&\theta_{m,k+1}\end{bmatrix},$
where, for all $j=1,\ldots,m$, $\theta_{j,k+1}\in{\mathbb{R}}^{n}$ is given by
$P_{k+1}^{-1}$ $=P_{k}^{-1}+\phi_{k}^{\rm T}\Gamma_{k}\phi_{k},$
$\theta_{j,k+1}$ $=\theta_{j,k}+P_{k+1}\phi_{k}^{\rm T}\Gamma_{k}(y_{j,k}-\phi_{k}\theta_{j,k}).$
Applying (20) and (22) then implies (37) and (38). ∎
Note that in independent identical column weighting, the same covariance matrix is used for each column of $\theta_{k}$ versus independent column weighting which computes a separate covariance matrix for each column of parameters $\theta_{j,k}$, $j=1,\ldots,m$. The computational complexities of BLS and RLS with independent identical column weighting are shown in Tables I and II respectively. Independent column weighting results in a $\mathcal{O}(m^{3})$ times improvement in computational complexity over vec-permutation as well as a $\mathcal{O}(m^{2})$ times improvement in space complexity over vec-permutation.
Figure 1 shows numerical testing of BLS and RLS with vec-permutation, independent column weighting, and independent identical column weighting. We consider the measurement process (1) with $p=10$, $n=50$, and $1\leq m\leq 20$. For larger values of $m$, we see significantly faster computation time for independent identical column weighting over vec-permutation and independent column weighting.
IV-A Convergence of Matrix RLS
It is well-known that in standard RLS, the parameter estimate vector converges to the vector of true parameters if the sequence of regressors $(\phi_{k})_{k=0}^{\infty}$ is persistently exciting [17, 18, 19]. Theorem 1 extends this result to matrix RLS. The following definition is from page 64 of [1].
@@ IV-A Convergence of Matrix RLS
$=(R+\sum_{i=0}^{k-1}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i})^{-1}(R\theta_{0}-R\theta+R\theta+\sum_{i=0}^{k-1}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i}\theta)$
@@
TABLE I: Batch identification summary and computational complexities for $N$ measurements
@@
Indep. Column Weight (23), (24), (25) $\mathcal{O}(\max\{Npnm\max\{p,n\},n^{3}m\})$ $\mathcal{O}(Npn^{2}m)$ $\mathcal{O}(Np^{2}nm)$
Indep. Iden. Column Weight (33), (34), (35) $\mathcal{O}(\max\{Np\max\{n,m\}\max\{p,n\},n^{3}\})$ $\mathcal{O}(Npn\max\{n,m\})$ $\mathcal{O}(Np^{2}\max\{n,m\})$
TABLE II: Recursive identification summary and computational/space complexities per step
@@
Indep. Column Weight (29), (28) $\mathcal{O}(pn^{2}m)$ (27), (28) $\mathcal{O}(p^{2}nm)$ $n^{2}m+nm$
Indep. Iden. Column Weight (39), (38) $\mathcal{O}(pn\max\{n,m\})$ (37), (38) $\mathcal{O}(p^{2}\max\{n,m\})$ $n^{2}+nm$
@@
V Application to Online Identification for Indirect Adaptive Model Predictive Control
Consider a MIMO input-output system of the form
@@ V Application to Online Identification for Indirect Adaptive Model Predictive Control
Proof.
This result follows from Lemma A.3. For further details, see equations (8) through (20) of [4]. ∎
@@
Proof.
Note that, for all $k\geq 0$, $z_{k}(\hat{\theta})^{\rm T}z_{k}(\hat{\theta})=\operatorname{tr}(z_{k}(\hat{\theta})^{\rm T}z_{k}(\hat{\theta}))=\operatorname{tr}(z_{k}(\hat{\theta})z_{k}(\hat{\theta})^{\rm T})$. It then follows from (57) of Lemma A.2 that, for all $k\geq 0$, (43) can be written as
$J_{k}(\hat{\theta})=\operatorname{tr}\Big{[}\sum_{i=0}^{k}z_{i}(\hat{\theta})z_{i}(\hat{\theta})^{\rm T}+(\hat{\theta}-\theta_{0}){P}_{0}^{-1}(\hat{\theta}-\theta_{0})^{\rm T}\Big{]}.$
Proposition 5 then implies that, for all $k\geq 0$, (53) and $\theta_{k+1}^{\rm T}=\theta_{k}^{\rm T}+P_{k+1}\phi_{k}(y_{k}^{\rm T}-\phi_{k}^{\rm T}\theta_{k}^{\rm T})$ hold. Taking the transpose then yields (54). ∎
Example 1. This example is from [10] and uses PCAC for the control of a flexible structure under harmonic and broadband disturbances. Consider the 4-bay truss show in Figure 2 made of flexible truss elements with unknown mass and stiffness. Two actuators are placed at nodes 3 and 4 with control authority in the $x$-direction and $x$-direction displacement sensors are placed at nodes 5, 6, 7, and 8. The objective is to use PCAC to suppress the effects of harmonic and broadband disturbances without prior knowledge of the truss dynamics. See [10] for further details.
@@
TABLE III: Truss ex. computation time per step: mean $\pm$ std. deviation
@@
VI Conclusions
This work derives batch and recursive least squares algorithms for the identification of matrix parameters. Under the assumption of independent, identical column weighting, these methods minimize the same cost function as the vec-permutation approach. It is also shown how, under persistent excitation, convergence guarantees can be extended from the vector case to the matrix case. This approach can be used fast online identification of MIMO systems which is critical in indirect adaptive model predictive control.

### v2

@@ 4 Matrix Update Least Squares
Propositions 3 show that, under assumptions (29) and (30), the cost function $J_{k}$, given by (4), can be minimized as a single matrix equation. We call this the matrix update approach.
@@
For all $k\geq 0$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let ${\Gamma}_{k}\in{\mathbb{R}}^{p\times p}$ be positive definite. Furthermore, let ${\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let ${R}\in{\mathbb{R}}^{n\times n}$ be positive definite. Then, for all $k\geq 0$, $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (4), has a unique minimizer, denoted as $\theta_{k+1}\triangleq\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta})$, which is given by
$\theta_{k+1}=A_{k}^{-1}b_{k},$ (32)
@@
$A_{k}$ $\triangleq R+\sum_{i=0}^{k}\phi_{i}^{\rm T}\Gamma_{i}\phi_{i},$ (33)
$b_{k}$ $\triangleq R\theta_{0}+\sum_{i=0}^{k}\phi_{i}^{\rm T}\Gamma_{i}y_{i}.$ (34)
Furthermore, for all $k\geq 0$, $\theta_{k+1}\in{\mathbb{R}}^{n\times m}$ is given recursively by
@@
where $P_{0}\triangleq R^{-1}$ and, for all $k\geq 0$, $P_{k}\in{\mathbb{R}}^{n\times n}$ is positive definite, hence nonsingular. Moreover, for all $k\geq 0$, $P_{k+1}$ can be expressed recursively as
@@
Note that (4) can be rewritten as (17) where, for all $k\geq 0$ and $j=1,\ldots,m$, $\Gamma_{j,k}=\Gamma_{k}$ and $R_{j}=R$. It then follows from Proposition 2 that $J_{k}$ has a unique minimizer given by $\theta_{k+1}=\begin{bmatrix}A_{k}^{-1}b_{1,k}&\cdots&A_{k}^{-1}b_{m,k}\end{bmatrix}=A_{k}^{-1}\begin{bmatrix}b_{1,k}&\cdots&b_{m,k}\end{bmatrix}$, where $A_{k}\in{\mathbb{R}}^{n\times n}$ is defined in (33) and $b_{1,k},\ldots,b_{m,k}$ are defined in (25). Finally, note that $b_{k}=[b_{1,k}\ \cdots\ b_{m,k}]$, yielding (32). Moreover, it also follows from Proposition 2 that, for all $k\geq 0$, $P_{k+1}^{-1}=P_{k}^{-1}+\phi_{k}^{\rm T}\Gamma_{k}\phi_{k},$ and $\theta_{j,k+1}=\theta_{j,k}+P_{k+1}\phi_{k}^{\rm T}\Gamma_{k}(y_{j,k}-\phi_{k}\theta_{j,k})$, where, for all $j=1,\ldots,m$, $\theta_{j,k}\in{\mathbb{R}}^{n}$ and $y_{j,k}\in{\mathbb{R}}^{p}$ are the $j^{\rm th}$ columns of $\theta_{k}$ and $y_{k}$, respectively. Combining columns yields (36) and applying matrix inversion lemma to (35) yields (37). ∎
Similarly, while the matrix update approach can be derived by applying standard least squares methods to the $m$ columns of (1) and combining columns [2, p. 103], our derivation shows how the matrix update approach implicitly implies independent and identical weighting to the columns of the residual error and parameter regularization.
4.1 Convergence of Matrix Update RLS
It is well-known that in standard RLS, the parameter estimate vector converges to the vector of true parameters if the sequence of regressors $(\phi_{k})_{k=0}^{\infty}$ is persistently exciting [16, 17]. Theorem 1 extends this result to matrix RLS. To begin, we extend the definition of persistent excitation (PE) from page 64 of [1] to the case of matrix regressors and nonuniform weight.222 Note that [1] considers a measurement process $y_{k}=\phi_{k}^{\rm T}\theta$ while we consider $y_{k}=\phi_{k}\theta$. As such, PE is defined in [1] using the limit $\lim_{k\rightarrow\infty}\frac{1}{k}\sum_{i=0}^{k-1}\phi_{i}\phi_{i}^{\rm T}$ whereas we consider $\lim_{k\rightarrow\infty}\frac{1}{k}\sum_{i=0}^{k-1}\phi_{i}^{\rm T}\phi_{i}$.
@@ 4.1 Convergence of Matrix Update RLS
Table 1: Batch least squares summary and computational complexities for $N$ measurements
@@
Column-by-Column (23), (24), (25) $\mathcal{O}(\max\{Npnm\max\{p,n\},n^{3}m\})$ $\mathcal{O}(Npn^{2}m)$ $\mathcal{O}(Np^{2}nm)$
Matrix Update (32), (33), (34) $\mathcal{O}(\max\{Np\max\{n,m\}\max\{p,n\},n^{3}\})$ $\mathcal{O}(Npn\max\{n,m\})$ $\mathcal{O}(Np^{2}\max\{n,m\})$
Table 2: Recursive least squares summary and computational/space complexities per step
@@
Column-by-Column (28), (27) $\mathcal{O}(pn^{2}m)$ (26), (27) $\mathcal{O}(p^{2}nm)$ $n^{2}m+nm$
Matrix Update (37), (36) $\mathcal{O}(pn\max\{n,m\})$ (35), (36) $\mathcal{O}(p^{2}\max\{n,m\})$ $n^{2}+nm$
@@
5 Computational Complexity and Performance Tradeoff
Next, we study the computational complexities of vec-permutation, column-by-column, and matrix update least squares. For simplicity, we assume it takes ${\mathcal{O}}(nmp)$ arithmetic operations to multiply an $(n\times m)$ matrix with a $(m\times p)$ matrix, ${\mathcal{O}}(n^{3})$ arithmetic operations to invert an $(n\times n)$ matrix, and ${\mathcal{O}}(nm)$ arithmetic operations to add two $(n\times m)$ matrices.
The computational complexities of the batch and recursive least squares methods are shown in Tables 1 and 2 respectively. Note that for the RLS methods, columns 2 and 4 of Table 2 show the most efficient implementation depending on the dimensions $n$ and $p$. Note that for both BLS and RLS, column-by-column and matrix update result in an $\mathcal{O}(m^{2})$ and $\mathcal{O}(m^{3})$ times improvement in computational complexity over vec-permutation, respectively. Moreover, for RLS, column-by-column and matrix update result in an $\mathcal{O}(m)$ and $\mathcal{O}(m^{2})$ times improvement in space complexity over vec-permutation, respectively. Next, Figure 1 shows numerical testing of BLS and RLS with vec-permutation, column-by-column, and matrix update. We consider the measurement process (1) with $p=10$, $n=50$, and $1\leq m\leq 20$. For larger values of $m$, we see significantly faster computation time for matrix update over vec-permutation and column-by-column.
While matrix update least squares offers a significant improvement in computational cost for large values of $m$, the following example shows that there may be a sacrifice in performance if there is prior knowledge that measurement noise has highly correlated columns. While not shown in this example, there may also be performance sacrifice if there is prior knowledge that columns of the parameters are highly correlated.
Example 5.1.
Consider the measurement process (1) with $p=2$, $m=2$, and $n=100$. We consider $10$ independent trials where, for each trial, the two columns of the parameters $\theta$ are i.i.d. sampled from the Gaussian distribution $\mathcal{N}(0_{n\times 1},I_{n})$. For each trial, for all $k\geq 0$, the two rows of the regressor, ${\phi}_{k}$ are i.i.d. sampled from $\mathcal{N}(0_{n\times 1},I_{n})$, and the measurement $y_{k}$ is given by $y_{k}=\phi_{k}\theta+v_{k}$, where $v_{k}\in{\mathbb{R}}^{p\times m}$ is the measurement noise and $\operatorname{vec}(v_{k})$ is i.i.d. sampled from $\mathcal{N}(0_{pm\times 1},\Sigma)$, where
[…]
and $\Sigma_{22}\triangleq 100\Sigma_{11}$. In this setup, the second column of the measurement noise has higher variance than that of the first column, and any two elements of the measurement noise are highly correlated with a correlation coefficient of $0.99$.
We compare the performance of vec-permutation, column-by-column, and matrix update RLS. We set regularization terms as $\bar{\theta}_{0}=0_{nm\times 1}$ and $\bar{R}=I_{nm}$, $\theta_{j,0}=0_{n\times 1}$ and $R_{j}=I_{n}$ $j=1,2$, and $\theta_{0}=0_{n\times m}$ and $R=I_{n}$, respectively. For vec-permutation, we let $\bar{\Gamma}_{k}=\Sigma^{-1}$, giving the minimum variance estimator. For column-by-column, we let $\Gamma_{1,k}=\Sigma_{11}^{-1}$ and $\Gamma_{2,k}=\Sigma_{22}^{-1}$. Finally, for matrix update, we consider the three choices $\Gamma_{k}=I_{2}$, $\Gamma_{k}=\Sigma_{11}^{-1}$, and $\Gamma_{k}=\Sigma_{22}^{-1}$.
[…]
6 Application to Online Identification for Indirect Adaptive Model Predictive Control
Finally, a useful application of this work is efficient online identification for adaptive model predictive control. Consider a MIMO input-output system of the form
@@ 6 Application to Online Identification for Indirect Adaptive Model Predictive Control
Proof 6.1.
This result follows from Proposition 1. For further details, see equations (8) through (20) of [4].
@@
Proof 6.2.
This result follows from Proposition 3.
Example 6.3.
This example is from [10] and uses PCAC for the control of a flexible structure under harmonic and broadband disturbances. Consider the 4-bay truss show in Figure 3 made of flexible truss elements with unknown mass and stiffness. Two actuators are placed at nodes 3 and 4 with control authority in the $x$-direction and $x$-direction displacement sensors are placed at nodes 5, 6, 7, and 8. The objective is to use PCAC to suppress the effects of broadband disturbances with constraints on actuator force and without prior knowledge of the truss dynamics. See [10] for further details.
@@
Table 3: Truss ex. computation time per step: mean $\pm$ std. deviation
@@
7 Conclusions
This work derives batch and recursive least squares algorithms for the identification of matrix parameters. Under the assumption of independent, identical column weighting, this method minimizes the same cost function as the vec-permutation approach while significantly improving computational complexity. It is also shown how, under persistent excitation, convergence guarantees can be extended from the vector case to the matrix case. This approach can be used fast online identification of MIMO systems which is critical in indirect adaptive model predictive control. A future area of interest is studying how various RLS forgetting algorithms (e.g. [19]) can be applied to matrix update RLS.